OpenStudy (anonymous):

Find the inverse formula for: f(x) = (2x-3)/(x+1)

OpenStudy (anonymous):

\[f(x) = y = \dfrac{2x-3}{x+1}\]\[f^{-1}(x) = x = \dfrac{2y - 3}{ y + 1}\] Now, find the y in f⁻¹(x)

OpenStudy (anonymous):

The answer is \[-\dfrac{x+3}{x-2}\]